# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@ 7. Toda lattice
Similarly, we denote by $\mathrm{HC}_{\hbar}(\mathrm{GL}_{N}^{L})$ the Harish-Chandra category $\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$ such that the $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$-bimodule structure on $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ is induced from the left action. For any $V\in\mathrm{Rep}(\mathrm{GL}_{N})$, define
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes V\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N}^{R}),\qquad\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{L})\otimes V\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N}^{L})$
the corresponding free Harish-Chandra bimodules of Section˜3.
For the rest of the section, we fix $\psi=1$ in (5.1). For any $X\in\mathrm{HC}_{\hbar}(\mathrm{GL}^{L}_{N})$, denote by
@@
the Kostant-Whittaker reduction of Section˜5 for left Harish-Chandra bimodules. In particular,
$N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N}^{R}).$
As in ˜4.2, for $x\in X$, we denote by
@@
correspondingly the image of $x$ under the projection map $X\rightarrow X/\mathfrak{n}_{+}^{L,-\psi}$ and the action of the Kirillov projector from Theorem˜5.2 on the latter. Recall that if $xP_{-\psi}^{L}$ is $N^{L}_{+}$-invariant, then we have $P_{-\psi}^{L}xP_{-\psi}^{L}=xP_{-\psi}^{L}$.
Likewise, for any $X\in\mathrm{HC}_{\hbar}(\mathrm{GL}^{R}_{N})$, denote by
$X/\!\!/N_{-}^{R}:=(\mathfrak{n}_{-}^{R}\backslash X)^{N_{-}^{R}},\qquad X/\!\!/_{\psi}\,N^{R}_{+}:=(X/\mathfrak{n}_{+}^{R,\psi})^{N_{+}^{R}}$
the parabolic restriction of Section˜4 and the Kostant-Whittaker reduction respectively for right Harish-Chandra bimodules. In particular, both
@@
are objects of $\mathrm{HC}_{\hbar}(\mathrm{GL}_{N}^{L})$. Similarly to the left action, we use notations
@@
for the extremal projector as in ˜4.2.
@@
Under the embeddings of Example˜3.5, we have
${}^{R}\iota_{\alpha}(\Omega(v_{i}))$ $=(E^{R}X)_{i\alpha},\qquad$ ${}^{R}\iota_{\alpha}^{*}(\Omega^{*}(\phi_{i}))$ $=(X^{-1}E^{R})_{\alpha i},$
${}^{L}\iota_{\alpha}^{*}(\Omega^{*}(\phi_{i}))$ $=(XE^{L})_{\alpha i},\qquad$ ${}^{L}\iota_{\alpha}(\Omega(v_{i}))$ $=(E^{L}X^{-1})_{i\alpha}.$
Since $v_{1}\in\mathbf{C}^{N}$ (resp. $\phi_{N}\in(\mathbf{C}^{N})^{*}$) is a highest-weight vector, the embeddings of Example˜3.5 factor through the maps of $\mathrm{GL}_{N}^{R}$ Harish-Chandra bimodules
(7.1) $\begin{split}{}^{R}\iota_{N}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})}{}^{R}\iota^{*}_{1}\colon\mathrm{U}_{\hbar}(\mathfrak{gl}^{R}_{N})\ltimes(S^{\bullet}\mathbf{C}^{N}\otimes S^{\bullet}(\mathbf{C}^{N})^{*})\rightarrow N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N}),\\ v_{i}\mapsto X_{iN},\qquad\phi_{i}\mapsto(X^{-1})_{1i},\end{split}$
The maps (7.1) descend to homomorphisms between corresponding reductions of Section˜4 and Section˜5. Denote by
(7.2) $Z_{iN}:={}^{R}\iota_{N}(\bar{v}_{i}),\qquad Z^{*}_{1i}:={}^{R}\iota^{*}_{1}(\bar{\phi}_{i}),\qquad Z_{iN},Z^{*}_{1j}\in N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}^{\mathrm{gen}}(\mathrm{GL}_{N})/\!\!/N^{R}_{-},$
in the notations of (4.10), (B.4), and (B.4), where $\mathrm{D}_{\hbar}^{\mathrm{gen}}(\mathrm{GL}_{N}):=\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})^{\mathrm{gen}}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})}\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$. Observe that the double reduction is an algebra.
@@
Using (7.1) and (7.2), it is enough to prove that $\bar{\phi}_{i}\bar{v}_{j}=\bar{v}_{j}\bar{\phi}_{i}$. This is the content of Proposition˜A.3. ∎
@@
We prove only the first equality. Recall the quantum comatrix $\hat{T}^{L}(u)$ from Section˜1.3 associated to vector fields generating left translations. Since they commute with the ones generating right translations, its matrix entries descend to elements of $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/N_{-}^{R}$, in other words, we have equality
$P^{R}\hat{T}^{L}_{1N}(u)=P^{R}\hat{T}^{L}_{1N}(u)P^{R}\in\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/N_{-}^{R},$
Using Proposition˜1.4 and Proposition˜A.4, we get
@@
In particular, we have $Z_{1a}^{*}Z_{aN}=P^{R}\hat{T}^{L}_{1N}(w_{a}^{R})P_{-\psi}^{L}$. At the same time, one can show analogously to [Kal25, Proposition 4.8] that
$\hat{T}_{1N}^{L}(u)P_{-\psi}^{L}=(-1)^{N+1}T_{2\ldots N}^{1\ldots N-1}(u)P_{-\psi}^{L}=\psi^{N+1}P_{-\psi}^{L}\in\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{L})/\mathfrak{n}_{+}^{L,-\psi}.$
@@
  1. (1)
    The algebra $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ is generated over $\mathbf{Z}_{\mathfrak{gl}_{N}}$ by $(E^{R})^{k}X]_{1N}P_{\psi}^{R}$ and $X^{-1}(E^{R})^{k}]_{1N}P_{\psi}^{R}$ for every $0\leq k\leq N-1$.
  2. (2)
    The algebra $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}^{\mathrm{gen}}/\!\!/N_{+}^{R}$ is generated by $Z_{iN}$ and $Z^{*}_{1i}$ from (7.2) over $\mathrm{U}_{\hbar}(\mathfrak{h}_{N}^{R})^{\mathrm{gen}}$ for $1\leq i\leq N$.
@@
$\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes\mathcal{O}(\mathrm{GL}_{N})\xrightarrow{\sim}N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{L})\ltimes\mathcal{O}(\mathrm{GL}_{N})\cong N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N}),\qquad f\mapsto P_{-\psi}^{L}fP_{-\psi}^{L}.$
In particular, we see that $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ is generated over $\mathbf{Z}_{\mathfrak{gl}_{N}}$ by $P_{-\psi}^{L}X_{ij}P_{-\psi}^{L}$ and $P_{-\psi}^{L}(X^{-1})_{ij}P_{-\psi}^{L}$ for all $i,j$. By Proposition˜5.7 and its dual version of Proposition˜B.2, it is equivalently generated by the classes of matrix entries $X(E^{L})^{k}]_{\alpha N}P_{-\psi}^{L}$ and $(E^{L})^{k}X^{-1}]_{1\alpha}P_{-\psi}^{L}$ in view of Proposition˜7.2, for every $\alpha$ and $k$. Therefore, by (1.14), $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ is generated by matrix entries $(E^{R})^{k}X]_{\alpha N}$ and $X^{-1}(E^{R})^{k}]_{1\alpha}$; for readability, we drop the sign $P_{-\psi}^{L}$.
@@
$P_{\psi}^{R}[(E^{R})^{k}X]_{\alpha N}P_{\psi}^{R},\qquad P_{\psi}^{R}[X^{-1}(E^{R})^{k}]_{1\alpha}P_{\psi}^{R}.$
@@
$P_{\psi}^{R}[(E^{R})^{k}X]_{\alpha N}P_{\psi}^{R}={}^{R}\iota_{N}(P_{\psi}\Omega^{k}(v_{i})P_{\psi}),\qquad P_{\psi}^{R}[X^{-1}(E^{R})^{k}]_{1\alpha}P_{\psi}^{R}={}^{R}\iota_{1}^{*}(P_{\psi}(\Omega^{*})^{k}(\phi_{i})).$
Therefore, by ˜5.4, its dual version of Proposition˜B.1, and Proposition˜3.7, the double Hamiltonian reduction is generated by the $N_{+}$-invariant classes
$^{R}\iota_{N}(\Omega^{k}(v_{1})P_{\psi})=[(E^{R})^{k}X]_{1N}P_{\psi}^{R},\qquad{}^{R}\iota_{1}^{*}((\Omega^{*})^{k}(\phi_{N})P_{\psi})=[X^{-1}(E^{R})^{k}]_{1N}P_{\psi}^{R}$
@@
For the second part, it follows that the double reduction $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}^{\mathrm{gen}}(\mathrm{GL}_{N})/\!\!/N_{-}^{R}$ is generated over $\mathrm{U}_{\hbar}(\mathfrak{h}_{N}^{R})^{\mathrm{gen}}$ by
$P^{R}[(E^{R})^{k}X]_{\alpha N}P^{R},\qquad P^{R}[X^{-1}(E^{R})^{k}]_{1\alpha}P^{R}.$
By ˜4.6 and its dual version (B.6), it is equivalently generated by (7.2), as required. ∎
Recall the algebra of difference operators $^{N}\tilde{\mathcal{A}}_{\hbar}$ from Section˜4.4. We will need the following result which is an analog of simplicity of Weyl algebras.
@@
It is enough to prove the lemma for the polynomial version $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})[\mathbf{u}_{1}^{\pm},\ldots,\mathbf{u}_{N}^{\pm}]$. Assume there is an ideal $I$ inside the latter and $a\in I$ is some element. We can write
$a=\sum_{J=(j_{1},\ldots,j_{N})}c_{J}(w)\mathbf{u}_{1}^{j_{1}}\ldots\mathbf{u}_{N}^{j_{N}},\qquad c_{J}(w)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N}).$
@@
$\mathbf{u}_{i},c(w_{1},\ldots,w_{N})]\mathbf{u}_{i}^{-1}=c(w_{1},\ldots,w_{i}+\hbar,\ldots,w_{N})-c(w_{1},\ldots,w_{i},\ldots,w_{N})$
@@
The $\mathrm{U}_{\hbar}(\mathfrak{h}_{N}^{R})^{\mathrm{gen}}$-linear map
@@
It is an algebra homomorphism by ˜7.3 and Proposition˜7.4. Surjectivity follows from the second part of ˜7.5. If it had a kernel, then it would be generated by some ideal in $\mathbf{C}[\hbar]$ by ˜7.6, in particular, the action of $\mathbf{C}[\hbar]$ on the class of $1]\in N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/N_{-}^{R}$ would have a kernel as well. Consider the Bruhat cell $U=B_{+}N_{-}\subset\mathrm{GL}_{N}$, where $B_{+}$ is the Borel subgroup. Since $(N_{+}^{L},N^{R}_{-})$-action is free, we have
@@
Finally, recall the finite Miura transform of ˜6.4:
(7.3) $\mathrm{Miura}_{N}\colon N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N^{R}_{+}\rightarrow N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}^{\mathrm{gen}}(\mathrm{GL}_{N})/\!\!/N^{R}_{-}.$
@@
Since $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})\cong\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes\mathcal{O}(\mathrm{GL}_{N})$ is a free module, it follows from trivializations (4.8) and (5.3) that this map is injective. By exactness of the Kostant-Whittaker reduction [BF08, Lemma 4(a)], the induced map
@@
The following theorem is an explicit presentation of the isomorphism from [BFN19, Appendix B].
@@
$x_{+}(u)$ $\mapsto-[(u-E^{R})^{-1}X]_{1N},\qquad$ $x_{-}(u)$ $\mapsto[X^{-1}(u-E^{R})^{-1}]_{1N},$

### v4

@@ 8 Toda lattice
Similarly, we denote by $\mathrm{HC}_{\hbar}\bigl(\mathrm{GL}_{N}^{L}\bigr)$ the Harish-Chandra category $\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$ such that the $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$-bimodule structure on $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ is induced from the left action. For any $V\!\in\!\mathrm{Rep}(\mathrm{GL}_{N})$, define
$\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{R}\bigr)\otimes V\in\mathrm{HC}_{\hbar}\bigl(\mathrm{GL}_{N}^{R}\bigr),\qquad\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{L}\bigr)\otimes V\in\mathrm{HC}_{\hbar}\bigl(\mathrm{GL}_{N}^{L}\bigr)$
the corresponding free Harish-Chandra bimodules of Section 4.
For the rest of the section, we fix $\psi=1$ in (6.1). For any $X\in\mathrm{HC}_{\hbar}\bigl(\mathrm{GL}^{L}_{N}\bigr)$, denote by
@@
the Kostant–Whittaker reduction of Section 6 for left Harish-Chandra bimodules. In particular,
$N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})\in\mathrm{HC}_{\hbar}\bigl(\mathrm{GL}_{N}^{R}\bigr).$
As in Notation 5.2, for $x\in X$, we denote by
@@
correspondingly the image of $x$ under the projection map $X\rightarrow X/\mathfrak{n}_{+}^{L,-\psi}$ and the action of the Kirillov projector from (6.2) on the latter. Recall that if $xP_{-\psi}^{L}$ is $N^{L}_{+}$-invariant, then we have $P_{-\psi}^{L}xP_{-\psi}^{L}=xP_{-\psi}^{L}$.
Likewise, for any $X\in\mathrm{HC}_{\hbar}\bigl(\mathrm{GL}^{R}_{N}\bigr)$, denote by
$X/\!\!/N_{-}^{R}:=\bigl(\mathfrak{n}_{-}^{R}\backslash X\bigr)^{N_{-}^{R}},\qquad X/\!\!/_{\psi}\,N^{R}_{+}:=\bigl(X/\mathfrak{n}_{+}^{R,\psi}\bigr)^{N_{+}^{R}}$
the parabolic restriction of Section 5 and the Kostant–Whittaker reduction respectively for right Harish-Chandra bimodules. In particular, both
@@
are objects of $\mathrm{HC}_{\hbar}\bigl(\mathrm{GL}_{N}^{L}\bigr)$. Similarly to the left action, we use notations
@@
for the extremal projector as in Notation 5.2.
In these notations, the quantum Toda lattice is $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N^{R}_{+}$.
Remark 8.2.
For simplicity, let $\hbar=1$. Character $\psi\colon\mathfrak{n}_{+}\rightarrow\mathbf{C}$ gives a multiplicative character of $N_{+}$ that we denote by ${\rm e}^{\psi}\colon N_{+}\rightarrow\mathbf{C}^{*}$. Let $U\subset\mathrm{GL}_{N}$ be the big Bruhat cell $U=N_{+}w_{0}H_{N}N_{+}$, where $w_{0}$ is the longest element in the Weyl group. Denote by
$C^{\infty}(U)\supset C^{\infty}_{(N^{L}_{+}\times N_{+}^{R},-\psi\times\psi)}(U):=\bigl\{f(n_{1}gn_{2})={\rm e}^{-\psi}(n_{1}){\rm e}^{\psi}(n_{2})f(g),\,n_{1},n_{2}\in N_{+}\bigr\}$ (8.1)
the space of smooth functions on $U$ equivariant with respect to $N_{+}^{L}\times N_{+}^{R}$-action and character $\bigl({\rm e}^{-\psi},{\rm e}^{\psi}\bigr)$. It can be identified with $C^{\infty}(H_{N})$, smooth functions on the torus.
There is a well-defined action of $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N^{R}_{+}$ on $C\raisebox{1.5pt}{${}^{\infty}_{(N^{L}_{+}\times N_{+}^{R},-\psi\times\psi)}$}(U)\cong C^{\infty}(H_{N})$. The quantum Toda Hamiltonians can be obtained as the images of the embedding
$Z(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}))\rightarrow N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N^{R}_{+}$
as bi-invariant differential operators twisted by character ${\rm e}^{\rho_{N}}\colon H_{N}\rightarrow\mathbf{C}^{*}$ (2.3). For instance, consider the operator
$C^{R}=\sum_{i=1}^{N}\bigl(E^{R}_{ii}\bigr)^{2}+2\sum_{i<j}E^{R}_{ji}E^{R}_{ij}+2\bigl(\rho_{N}^{\vee}\bigr)^{R}=\sum_{i,j=1}^{N}E^{R}_{ij}E^{R}_{ji}+(N-1)\sum_{i=1}^{N}E^{R}_{ii},$
a slight modification of the Casimir operator, where
$\rho_{N}^{\vee}=\sum_{i=1}^{N}(N-i)E_{ii}.$
Using transformation properties (8.1), one can check that
${\rm e}^{\rho_{N}}C^{R}{\rm e}^{-\rho_{N}}=\sum_{i=1}^{N}\partial_{i}^{2}-2\psi^{2}\sum_{i=1}^{N-1}{\rm e}^{x_{i}-x_{i+1}}-(\rho_{N},\rho_{N}),$
the standard Toda Hamiltonian. Here $\{{\rm e}^{x_{i}}\}$ are natural coordinates of diagonal entries on $H_{N}$.
@@
Under the embeddings of Example 4.5, we have
${}^{R}\iota_{\alpha}(\Omega(v_{i}))=\bigl(E^{R}X\bigr)_{i\alpha},\qquad$ ${}^{R}\iota_{\alpha}^{*}(\Omega^{*}(\phi_{i}))=\bigl(X^{-1}E^{R}\bigr)_{\alpha i},$
${}^{L}\iota_{\alpha}^{*}(\Omega^{*}(\phi_{i}))=\bigl(XE^{L}\bigr)_{\alpha i},\qquad$ ${}^{L}\iota_{\alpha}(\Omega(v_{i}))=\bigl(E^{L}X^{-1}\bigr)_{i\alpha}.$
Since $v_{1}\in\mathbf{C}^{N}$ (resp. $\phi_{N}\in\bigl(\mathbf{C}^{N}\bigr)^{*}$) is a highest-weight vector, the embeddings of Example 4.5 factor through the maps of $\mathrm{GL}_{N}^{R}$ Harish-Chandra bimodules
${}^{R}\iota_{N}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})}{}^{R}\iota^{*}_{1}\colon\ \mathrm{U}_{\hbar}\bigl(\mathfrak{gl}^{R}_{N}\bigr)\ltimes\bigl(S^{\bullet}\mathbf{C}^{N}\otimes S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)$ $\rightarrow N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N}),$
$v_{i}$ $\mapsto X_{iN},\qquad\phi_{i}\mapsto\bigl(X^{-1}\bigr)_{1i}.$ (8.2)
The maps (8.2) descend to homomorphisms between corresponding reductions of Section 5 and Section 6. Denote by
$Z_{iN}:={}^{R}\iota_{N}(\bar{v}_{i}),\qquad Z^{*}_{1i}:={}^{R}\iota^{*}_{1}\bigl(\bar{\phi}_{i}\bigr),\qquad Z_{iN},Z^{*}_{1j}\in N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}^{\mathrm{gen}}(\mathrm{GL}_{N})/\!\!/N^{R}_{-},$ (8.3)
in the notations of (5.9), (B.3), and (B.3), where $\mathrm{D}_{\hbar}^{\mathrm{gen}}(\mathrm{GL}_{N}):=\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{R}\bigr)^{\mathrm{gen}}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})}\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$. Observe that the double reduction is an algebra.
@@
Using (8.2) and (8.3), it is enough to prove that $\bar{\phi}_{i}\bar{v}_{j}=\bar{v}_{j}\bar{\phi}_{i}$. This is the content of Proposition A.3. ∎
@@
We prove only the first equality. Recall the quantum comatrix $\hat{T}^{L}(u)$ from Section 2.3 associated to vector fields generating left translations. Since they commute with the ones generating right translations, its matrix entries descend to elements of $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/N_{-}^{R}$, in other words, we have equality
$P^{R}\hat{T}^{L}_{1N}(u)=P^{R}\hat{T}^{L}_{1N}(u)P^{R}\in\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/N_{-}^{R}.$
Using Propositions 2.4 and A.4, we get
@@
In particular, we have $Z_{1a}^{*}Z_{aN}=P^{R}\hat{T}^{L}_{1N}\bigl(w_{a}^{R}\bigr)P_{-\psi}^{L}$. At the same time, one can show analogously to [27, Proposition 4.8] that
$\hat{T}_{1N}^{L}(u)P_{-\psi}^{L}=(-1)^{N+1}T_{2\dots N}^{1\dots N-1}(u)P_{-\psi}^{L}=\psi^{N+1}P_{-\psi}^{L}\in\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{L}\bigr)/\mathfrak{n}_{+}^{L,-\psi}.$
@@
  1. $1.$
    The algebra $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ is generated over $\mathbf{Z}_{\mathfrak{gl}_{N}}$ by $\bigl[\bigl(E^{R}\bigr)^{k}X\bigr]_{1N}P_{\psi}^{R}$ and $\bigl[X^{-1}\bigl(E^{R}\bigr)^{k}\bigr]_{1N}P_{\psi}^{R}$ for every $0\leq k\leq N-1$.
  2. $2.$
    The algebra $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}^{\mathrm{gen}}/\!\!/N_{+}^{R}$ is generated by $Z_{iN}$ and $Z^{*}_{1i}$ from (8.3) over $\mathrm{U}_{\hbar}\bigl(\mathfrak{h}_{N}^{R}\bigr)^{\mathrm{gen}}$ for $1\leq i\leq N$.
@@
$\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes\mathcal{O}(\mathrm{GL}_{N})$ $\xrightarrow{\sim}N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{L}\bigr)\ltimes\mathcal{O}(\mathrm{GL}_{N})\cong N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N}),$
$f$ $\mapsto P_{-\psi}^{L}fP_{-\psi}^{L}.$
In particular, we see that $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ is generated over $\mathbf{Z}_{\mathfrak{gl}_{N}}$ by the classes $P_{-\psi}^{L}X_{ij}P_{-\psi}^{L}$ and $P_{-\psi}^{L}\bigl(X^{-1}\bigr)_{ij}P_{-\psi}^{L}$ for all $i$, $j$. By Proposition 6.7 and its dual version of Proposition B.2, it is equivalently generated by the classes of matrix entries $\bigl[X\bigl(E^{L}\bigr)^{k}\bigr]_{\alpha N}P_{-\psi}^{L}$ and $\bigl[\bigl(E^{L}\bigr)^{k}X^{-1}\bigr]_{1\alpha}P_{-\psi}^{L}$ in view of Proposition 8.3, for every $\alpha$ and $k$. Therefore, by (2.11), $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ is generated by matrix entries $\bigl[\bigl(E^{R}\bigr)^{k}X\bigr]_{\alpha N}$ and $\bigl[X^{-1}\bigl(E^{R}\bigr)^{k}\bigr]_{1\alpha}$; for readability, we drop the sign $P_{-\psi}^{L}$.
@@
$P_{\psi}^{R}\bigl[\bigl(E^{R}\bigr)^{k}X\bigr]_{\alpha N}P_{\psi}^{R},\qquad P_{\psi}^{R}\bigl[X^{-1}\bigl(E^{R}\bigr)^{k}\bigr]_{1\alpha}P_{\psi}^{R}.$
@@
$P_{\psi}^{R}\bigl[\bigl(E^{R}\bigr)^{k}X\bigr]_{\alpha N}P_{\psi}^{R}={}^{R}\iota_{N}\bigl(P_{\psi}\Omega^{k}(v_{i})P_{\psi}\bigr),\qquad P_{\psi}^{R}\bigl[X^{-1}\bigl(E^{R}\bigr)^{k}\bigr]_{1\alpha}P_{\psi}^{R}={}^{R}\iota_{1}^{*}\bigl(P_{\psi}(\Omega^{*})^{k}(\phi_{i})\bigr).$
Therefore, by Lemma 6.4, its dual version of Proposition B.1, and Proposition 4.7, the double Hamiltonian reduction is generated by the $N_{+}$-invariant classes
$^{R}\iota_{N}\bigl(\Omega^{k}(v_{1})P_{\psi}\bigr)=\bigl[\bigl(E^{R}\bigr)^{k}X\bigr]_{1N}P_{\psi}^{R},\qquad{}^{R}\iota_{1}^{*}\bigl((\Omega^{*})^{k}(\phi_{N})P_{\psi}\bigr)=\bigl[X^{-1}\bigl(E^{R}\bigr)^{k}\bigr]_{1N}P_{\psi}^{R}$
@@
For the second part, it follows that the double reduction $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}^{\mathrm{gen}}(\mathrm{GL}_{N})/\!\!/N_{-}^{R}$ is generated over $\mathrm{U}_{\hbar}\bigl(\mathfrak{h}_{N}^{R}\bigr)^{\mathrm{gen}}$ by
$P^{R}\bigl[\bigl(E^{R}\bigr)^{k}X\bigr]_{\alpha N}P^{R},\qquad P^{R}\bigl[X^{-1}\bigl(E^{R}\bigr)^{k}\bigr]_{1\alpha}P^{R}.$
By Lemma 5.6 and its dual version (B.5), it is equivalently generated by (8.3), as required. ∎
Recall the algebra of difference operators $^{N}\!\tilde{\mathcal{A}}_{\hbar}$ from Section 5.4. We will need the following result which is an analog of simplicity of Weyl algebras.
@@
It is enough to prove the lemma for the polynomial version $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})\bigl[\mathbf{u}_{1}^{\pm},\dots,\mathbf{u}_{N}^{\pm}\bigr]$. Assume there is an ideal $I$ inside the latter and $a\in I$ is some element. We can write
$a=\sum_{J=(j_{1},\dots,j_{N})}c_{J}(w)\mathbf{u}_{1}^{j_{1}}\cdots\mathbf{u}_{N}^{j_{N}},\qquad c_{J}(w)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N}).$
@@
$\mathbf{u}_{i},c(w_{1},\dots,w_{N})]\mathbf{u}_{i}^{-1}=c(w_{1},\dots,w_{i}+\hbar,\dots,w_{N})-c(w_{1},\dots,w_{i},\dots,w_{N}).$
@@
The $\mathrm{U}_{\hbar}\bigl(\mathfrak{h}_{N}^{R}\bigr)^{\mathrm{gen}}$-linear map
@@
It is an algebra homomorphism by Lemma 8.4 and Proposition 8.5. Surjectivity follows from the second part of Lemma 8.6. If it had a kernel, then it would be generated by some ideal in $\mathbf{C}[\hbar]$ by Lemma 8.7, in particular, the action of $\mathbf{C}[\hbar]$ on the class of $1]\!\in\!N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/N_{-}^{R}$ would have a kernel as well. Consider the Bruhat cell $U=B_{+}N_{-}\subset\mathrm{GL}_{N}$, where $B_{+}$ is the Borel subgroup. Since $\bigl(N_{+}^{L},N^{R}_{-}\bigr)$-action is free, we have
@@
Finally, recall the finite Miura transform of Definition 7.4:
$\mathrm{Miura}_{N}\colon\ N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N^{R}_{+}\rightarrow N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}^{\mathrm{gen}}(\mathrm{GL}_{N})/\!\!/N^{R}_{-}.$ (8.4)
@@
Since $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})\cong\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{R}\bigr)\otimes\mathcal{O}(\mathrm{GL}_{N})$ is a free module, it follows from trivializations (5.7) and (6.3) that this map is injective. By exactness of the Kostant–Whittaker reduction [3, Lemma 4 (a)], the induced map
@@
The following theorem is an explicit presentation of the isomorphism from [4, Appendix B].
@@
$x_{+}(u)\mapsto-\bigl[\bigl(u-E^{R}\bigr)^{-1}X\bigr]_{1N},\qquad$ $x_{-}(u)\mapsto\bigl[X^{-1}\bigl(u-E^{R}\bigr)^{-1}\bigr]_{1N},$